# “It Explains What I am Currently Going Through Perfectly to a Tee”: Understanding User Perceptions on LLM-Enhanced Narrative Interventions

### v1

@@ 4.1. LLM-Enhanced Stories in Exhibiting Key Narrative Qualities
Crowdworkers ($\textbf{n}_{Non-LLM}=\textbf{n}_{LLM}=87$) Students ($\textbf{n}_{Non-LLM}=90,\textbf{n}_{LLM}=82$)
Mean Mean
Quality Hypothesized Superior Group Non-LLM LLM p-value Non-LLM LLM p-value
Perceived Authenticity Non-LLM 5.03 $\pm$ 0.17 4.92 $\pm$ 0.17 0.31 4.34 $\pm$ 0.18 4.29 $\pm$ 0.19 0.42
Perceived Ability to Communicate Key Takeaways LLM 5.51 $\pm$ 0.14 6.01 $\pm$ 0.10 ¡0.01** 5.27 $\pm$ 0.16 5.48 $\pm$ 0.17 0.21
Perceived Ability to Promote Reflection 4.71 $\pm$ 0.20 5.31 $\pm$ 0.17 0.02* 3.96 $\pm$ 0.20 4.55 $\pm$ 0.20 0.04*
Perceived Balance of Positivity with Realistic Struggle 4.80 $\pm$ 0.16 5.38 $\pm$ 0.16 ¡0.01** 4.29 $\pm$ 0.18 4.62 $\pm$ 0.17 0.15
@@ 4.1.1. Perceived Authenticity
Participants in both LLM and Non-LLM conditions were drawn to the authenticity of their stories because of their relatability and the sincerity depicted in the characters’ journeys. These narratives were appreciated for their candid portrayal of often unspoken feelings, aligning closely with personal and broadly shared human experiences.
However, for some participants in the Non-LLM condition, the very universality that made the stories relatable also rendered them suspect in terms of authenticity. They perceived these universal themes as too generic, suggesting a lack of specificity that could make the stories feel impersonal.
Participants in the LLM condition provided various insights into the factors contributing to perceived authenticity in the LLM-enhanced stories. SL473 remarked on the relatability and credibility of their story, stating:
@@
The quality of the writing also played a crucial role in enhancing the authenticity of the stories. Descriptive and vivid prose helped participants clearly visualize the scenarios, fostering a deeper connection with the content. CL73 shared, “I could picture the scenario pretty well in my mind and resonate with it.” However, several participants criticized the tone and writing style of their LLM-enhanced stories as being too mechanical and impersonal. Doubts about AI involvement were raised due to perceived shortcomings in the depth of the protagonists and their emotions. In fact, some participants commented that their stories shared patterns similar to the typical responses made by ChatGPT. CL41 noted:
@@
Furthermore, some participants felt that the LLM-enhanced stories were overly tailored to their specific situations, making the AI’s involvement too apparent and diminishing the authenticity of the narratives. This precise mirroring of their inputs led to a perception that the stories were merely processing and regurgitating information, which reduced their impact and made them feel less genuine.
@@ 4.1.2. Perceived Ability to Communicate Key Takeaways
Participants in both conditions valued the structured progression within their stories, which clearly delineated the flow from beginning to end. However, some participants in the Non-LLM condition commented that the takeaways were not very evident as they did not relate to their stories. For instance, SN63 acknowledged understanding the message but felt a disconnect from their personal experiences.
@@
Criticisms of the LLM-enhanced stories in this regard were minimal, mainly focusing on the tone and presentation of the content. The feedback echoed previous sentiments, highlighting that mechanical and impersonal tones detracted from the authenticity of the narratives. Additionally, there was some concern from the students that advice that seemed to be generated by AI might not be taken as seriously by readers, potentially undermining the impact and credibility of the guidance provided.
@@ 4.1.3. Perceived Ability to Promote Reflection
Participants in the Non-LLM condition valued how their stories depicted common experiences that reinforced actions and introduced fresh perspectives, but they were sometimes criticized for being too generic. This generality made their stories seem less reliable as catalysts for deep reflection. The lack of specific, detailed scenarios in these stories sometimes led participants to perceive the stories as insufficiently engaging or thought-provoking.
@@ 4.1.4. Perceived Balance of Positivity with Realistic Struggles
The structured progression of events in the stories was viewed favorably in both conditions, as it helped maintain a balance between positivity and realistic portrayals of struggles, anchoring the narratives and ensuring they resonated with the audience to some extent. However, participants in the LLM condition were particularly appreciative of the balance achieved between positivity and realistic struggles They particularly highlighted their stories’ empathetic acknowledgment of the protagonist’s struggles, which mirrored the challenges faced by the readers themselves. This recognition of shared difficulties established a foundation of empathy, making participants feel understood and more connected to the narrative. Additionally, their stories were commended for illustrating the small, incremental steps protagonists took to address and improve their challenges, enhancing their realism and applicability.
Participants appreciated that their stories did not promise unattainable perfection but instead portrayed a more truthful journey towards improvement. This realistic depiction resonated well with participants like SL86, who stated:
@@
However, participants noted that both the Non-LLM and LLM-enhanced stories often adopted a more positive tone than they would have preferred. This sentiment was especially pronounced when participants perceived their situations as beyond their control, leading to a disconnect with the occasional overly optimistic narrative. CL35 expressed this concern:
@@
4.2. LLM-Enhanced Stories as an Intervention to Manage Negative Thoughts
Table 4 presents the summary statistics of the story attribute ratings given by participants. We observed that the stories from the LLM condition were perceived to be superior in all the qualities. Save for the ratings of perceived emotional intensity, the differences in ratings were statistically significant among the crowdworkers. Among the students, only the differences in perceived relatability, perceived capacity of inspiring solutions, and length of written reflection were statistically significant. Below, we supplement these findings with participant comments that highlight both the beneficial aspects and the areas for improvement of the stories.
@@ 4.2. LLM-Enhanced Stories as an Intervention to Manage Negative Thoughts
Crowdworkers ($\textbf{n}_{Non-LLM}=\textbf{n}_{LLM}=87$) Students ($\textbf{n}_{Non-LLM}=90,\textbf{n}_{LLM}=82$)
Mean Mean
Quality Hypothesized Superior Group Non-LLM LLM p-value Non-LLM LLM p-value
Perceived Reduction in Belief in Negative Thought LLM 0.26 $\pm$ 0.10 0.64 $\pm$ 0.13 0.02* 0.49 $\pm$ 0.12 0.65 $\pm$ 0.12 0.21
Perceived Reduction in Emotion Intensity 0.59 $\pm$ 0.13 0.89 $\pm$ 0.12 0.06 0.44 $\pm$ 0.10 0.65 $\pm$ 0.12 0.15
Perceived Relatability 5.16 $\pm$ 0.17 5.75 $\pm$ 0.13 ¡0.01** 4.48 $\pm$ 0.17 5.05 $\pm$ 0.18 0.03*
Perceived Capacity of Inspiring Solutions 4.74 $\pm$ 0.18 5.46 $\pm$ 0.16 ¡0.01** 4.14 $\pm$ 0.19 4.76 $\pm$ 0.18 0.04*
Likelihood of Future Engagement 4.53 $\pm$ 0.20 5.21 $\pm$ 0.17 ¡0.01** 3.92 $\pm$ 0.20 4.30 $\pm$ 0.20 0.15
Number of Words Written for Reflection 31.15 $\pm$ 2.28 44.22 $\pm$ 4.04 ¡0.01** 25.29 $\pm$ 2.42 33.56 $\pm$ 2.65 0.04*
@@ 4.2.1. Reevaluating Existing Thoughts and Beliefs
The reevaluation process also involved a reflective writing activity associated with the reflection question. As shown in Table 4, both crowdworkers and students wrote more words on average in response to the LLM-enhanced stories. This activity aimed to provide participants with an opportunity to organize their thoughts and emotions into written form, which may have helped them restructure their thoughts to be more helpful. SL28 expressed one such change, saying:
@@
These findings suggest that both Non-LLM and LLM-enhanced stories could support participants in reevaluating their thoughts and beliefs, offering them new insights and perspectives.
@@ 4.2.2. Potential of Stories in Inspiring Solutions and Actions
Participants expressed that this reflection process involved more than just passive reading. The LLM-enhanced stories helped participants to reflect on their personal struggles through the lens of the story and to actively engage with the suggested solutions. By presenting coping mechanisms embedded within the context of the narrative, participants felt the stories allowed them to see how similar strategies might apply in their own lives. The stories not only illustrated relatable challenges but also modeled potential solutions in a way that felt accessible and actionable. SL64 commented:
@@ 5. Discussion
Our study demonstrates how LLMs can significantly enhance human-generated content so that it is more relatable and impactful to new audiences. We found that these enhancements improved the utility of the stories as interventions for managing negative thoughts among young adults. These findings add to the growing body of literature on human-LLM collaboration (Chan et al., 2023; Denny et al., 2023; Bhattacharjee et al., 2024b; Guingrich and Graziano, 2023; Sharma et al., 2023a), which seeks to utilize and expand upon existing support mechanisms and content, while shining a light on the evaluation and scrutiny required to support the design of such tools for DMH (Bhattacharjee et al., 2024b; Korngiebel and Mooney, 2021; Demszky et al., 2023).
@@ 5.1. Key Insights
5.1.1. RQ1: LLM-enhanced Stories in Exhibiting Narrative Qualities
@@
Our research highlights the promising capabilities of LLMs to dynamically generate content that can enhance the quality of narratives in DMH interventions. Similar to prior works (Sharma et al., 2023b; Kim et al., 2024; Bhattacharjee et al., 2024b), the LLM-enhanced stories in our study processed open-ended inputs from participants and adapted the stories to their unique contexts. This capability underscores the potential of LLMs in crafting personalized stories that can support individuals in managing their specific barriers to psychological wellbeing.
@@
Our findings reveal that the LLM-enhanced stories in our study enhanced most of the narrative qualities we investigated. The LLM-enhanced stories were effective in communicating clear and actionable takeaways, which participants noted as essential for translating the stories’ lessons into their own lives (Bhattacharjee et al., 2022; Dyer, 1982). These stories were perceived to excel in promoting reflection, as explicit suggestions within the narratives encouraged readers to contemplate potential changes in their thinking patterns (Bhattacharjee et al., 2024a; Kocielnik et al., 2018). Additionally, the LLM-enhanced stories balanced positivity with realistic portrayals of struggles. The stories often charted a structured progression from challenging life scenarios that resonated with participants’ experiences to outcomes showing small yet meaningful progress (Bhattacharjee et al., 2022).
@@
The LLM was able to render all of these benefits while still being comparable in authenticity to human-written stories, which likely reflects that the LLM effectively mirrored participants’ life experiences and that it drew on human-written example stories. However, while not statistically significant, we also note that the trend in the difference between the two conditions, favored the Non-LLM conition. Our qualitative data also revealed concerns about authenticity in the LLM condition, which could potentially reflect specific word choices made by the LLMs, participants’ familiarity with typical responses made by popular chatbots (Ariyaratne et al., 2023), and the occasional mechanical tone inherent in AI-generated content (Cheng et al., 2022). Recent studies have also examined text generated by ChatGPT across various contexts, noting an increased and repetitive occurrence of certain words such as ‘delve’, ‘intricate’, and ‘underscore’ (Kobak et al., 2024). The frequent use of certain relatively formal words or phrases may contribute to compromising the perceived authenticity of LLM-generated stories.
@@
5.1.2. RQ2: LLM-Enhanced Stories as an Intervention to Help People Manage Negative Thoughts

### v4

@@ 4.1. Personalized LLM-Enhanced Stories in Exhibiting Key Narrative Qualities
[…]
Quantitative comparisons of perceived narrative qualities between LLM and Non-LLM stories, including authenticity, key takeaways, reflection, and balance of positivity with realistic struggles. LLM stories were rated significantly higher on several dimensions, particularly by crowdworkers. Crowdworkers ($\textbf{n}_{Non-LLM}=\textbf{n}_{LLM}=87$) Students ($\textbf{n}_{Non-LLM}=90,\textbf{n}_{LLM}=82$) Mean Mean Quality Hypothesized Superior Group Non-LLM LLM p-value Non-LLM LLM p-value Perceived Authenticity Non-LLM 5.03 $\pm$ 0.17 4.92 $\pm$ 0.17 0.31 4.34 $\pm$ 0.18 4.29 $\pm$ 0.19 0.42 Perceived Ability to Communicate Key Takeaways LLM 5.51 $\pm$ 0.14 6.01 $\pm$ 0.10 ¡0.01** 5.27 $\pm$ 0.16 5.48 $\pm$ 0.17 0.21 Perceived Ability to Promote Reflection 4.71 $\pm$ 0.20 5.31 $\pm$ 0.17 0.02* 3.96 $\pm$ 0.20 4.55 $\pm$ 0.20 0.04* Perceived Balance of Positivity with Realistic Struggle 4.80 $\pm$ 0.16 5.38 $\pm$ 0.16 ¡0.01** 4.29 $\pm$ 0.18 4.62 $\pm$ 0.17 0.15
@@ 4.1.1. Perceived Authenticity
Across both populations, perceived authenticity ratings were similar across conditions. Among crowdworkers, the LLM stories ($Mean_{\text{LLM}}$ = 4.92 $\pm$ 0.17) received slightly lower ratings than the Non-LLM stories ($Mean_{\text{Non-LLM}}$ = 5.03 $\pm$ 0.17), with no significant difference ($p$ = 0.31). Among students, ratings were nearly identical for both conditions ($Mean_{\text{LLM}}$ = 4.29 $\pm$ 0.19, $Mean_{\text{Non-LLM}}$ = 4.34 $\pm$ 0.18, $p$ = 0.42). Participants in both LLM and Non-LLM conditions were drawn to the authenticity of their stories because of their relatability and the sincerity depicted in the characters’ journeys. These narratives were appreciated for their candid portrayal of often unspoken feelings, aligning closely with personal and broadly shared human experiences. However, for some participants in the Non-LLM condition, the very universality that made the stories relatable also rendered them suspect in terms of authenticity. They perceived these universal themes as too generic, suggesting a lack of specificity that could make the stories feel impersonal.
Participants in the LLM condition provided various insights into the factors contributing to perceived authenticity in the LLM stories. SL473 remarked on the relatability and credibility of their story, stating:
@@
The quality of the writing also played a crucial role in the authenticity of the LLM stories. Descriptive and vivid prose helped participants clearly visualize the scenarios, fostering a deeper connection with the content. CL73 shared, “I could picture the scenario pretty well in my mind and resonate with it.” However, several participants criticized the tone and writing style of their stories as being too mechanical and impersonal. Doubts about AI involvement were raised due to perceived shortcomings in the depth of the protagonists and their emotions. In fact, some participants commented that their stories shared patterns similar to the typical responses made by ChatGPT. CL41 noted:
@@
Furthermore, some participants felt that the LLM stories were overly tailored to their specific situations, making the AI’s involvement too apparent and diminishing the authenticity of the narratives. This precise mirroring of their inputs led to a perception that the stories were merely processing and regurgitating information, which reduced their impact and made them feel less genuine.
@@ 4.1.2. Perceived Ability to Communicate Key Takeaways
LLM stories were rated significantly higher than Non-LLM stories in terms of communicating key takeaways among crowdworkers ($Mean_{\text{LLM}}$ = 6.01 $\pm$ 0.10 vs. $Mean_{\text{Non-LLM}}$ = 5.51 $\pm$ 0.14, $p$ 0.01). Among students, the LLM stories also received higher ratings ($Mean_{\text{LLM}}$ = 5.48 $\pm$ 0.17 vs. $Mean_{\text{Non-LLM}}$ = 5.27 $\pm$ 0.16), though the difference was not statistically significant ($p$ = 0.21). Participants in both conditions valued the structured progression within their stories, which clearly delineated the flow from beginning to end. However, some participants in the Non-LLM condition commented that the takeaways were not very evident as they did not relate to their stories. For instance, SN63 acknowledged understanding the message but felt a disconnect from their personal experiences.
@@
Criticisms of the LLM stories in this regard were minimal, mainly focusing on the tone and presentation of the content. The feedback echoed previous sentiments, highlighting that mechanical and impersonal tones detracted from the authenticity of the narratives. Additionally, there was some concern from the students that advice that seemed to be generated by AI might not be taken as seriously by readers, potentially undermining the impact and credibility of the guidance provided.
@@ 4.1.3. Perceived Ability to Promote Reflection
LLM stories were rated significantly higher in promoting reflection among both groups. For crowdworkers, the LLM condition received a mean rating of $Mean_{\text{LLM}}$ = 5.31 $\pm$ 0.17, compared to $Mean_{\text{Non-LLM}}$ = 4.71 $\pm$ 0.20 in the Non-LLM condition ($p$ = 0.02). Among students, the LLM stories also outperformed Non-LLM stories ($Mean_{\text{LLM}}$ = 4.55 $\pm$ 0.20 vs. $Mean_{\text{Non-LLM}}$ = 3.96 $\pm$ 0.20, $p$ = 0.04). Participants in the Non-LLM condition valued how their stories depicted common experiences that reinforced actions and introduced fresh perspectives, but they were sometimes criticized for being too generic. This generality made their stories seem less reliable as catalysts for deep reflection. The lack of specific, detailed scenarios in these stories sometimes led participants to perceive the stories as insufficiently engaging or thought-provoking.
@@ 4.1.4. Perceived Balance of Positivity with Realistic Struggles
Crowdworkers rated LLM stories significantly higher for balancing positivity with realistic struggles ($Mean_{\text{LLM}}$ = 5.38 $\pm$ 0.16) compared to Non-LLM stories ($Mean_{\text{Non-LLM}}$ = 4.80 $\pm$ 0.16, $p$ 0.01). Among students, the LLM condition also received higher ratings ($Mean_{\text{LLM}}$ = 4.62 $\pm$ 0.17 vs. $Mean_{\text{Non-LLM}}$ = 4.29 $\pm$ 0.18), though the difference was not statistically significant ($p$ = 0.15). The structured progression of events in the stories was viewed favorably in both conditions, as it helped maintain a balance between positivity and realistic portrayals of struggles, anchoring the narratives and ensuring they resonated with the audience to some extent. However, participants in the LLM condition were particularly appreciative of the balance achieved between positivity and realistic struggles They particularly highlighted their stories’ empathetic acknowledgment of the protagonist’s struggles, which mirrored the challenges faced by the readers themselves. This recognition of shared difficulties established a foundation of empathy, making participants feel understood and more connected to the narrative. Additionally, their stories were commended for illustrating the small, incremental steps protagonists took to address and improve their challenges, enhancing their realism and applicability.
Participants appreciated that their stories did not promise unattainable perfection but instead portrayed a more truthful journey toward improvement. This realistic depiction resonated well with participants like SL86, who stated:
@@
However, participants noted that both the Non-LLM and LLM stories often adopted a more positive tone than they would have preferred. This sentiment was especially pronounced when participants perceived their situations as beyond their control, leading to a disconnect with the occasional overly optimistic narrative. CL35 expressed this concern:
@@
4.2. Personalized LLM-Enhanced Stories as an Intervention to Manage Negative Thoughts
Table 4 presents the summary statistics of participant ratings on various intervention outcomes. We observed that the LLM stories were perceived to be superior in all the qualities. Additional details about ratings across different challenging thoughts are provided in Appendix B. Save for the ratings of perceived emotional intensity, the differences in ratings were statistically significant among the crowdworkers. Among the students, only the differences in perceived relatability, perceived capacity of inspiring solutions, and length of written reflection were statistically significant. Below, we supplement these findings with participant comments that highlight both the beneficial aspects and the areas for improvement of the stories.
@@ 4.2. Personalized LLM-Enhanced Stories as an Intervention to Manage Negative Thoughts
[…]
Quantitative comparisons of perceived intervention qualities, including reduction in belief in negative thoughts, emotion intensity, relatability, inspiration, likelihood of future engagement, and number of words written for reflection. LLM stories consistently outperformed Non-LLM stories, especially among crowdworkers. Crowdworkers ($\textbf{n}_{Non-LLM}=\textbf{n}_{LLM}=87$) Students ($\textbf{n}_{Non-LLM}=90,\textbf{n}_{LLM}=82$) Mean Mean Quality Hypothesized Superior Group Non-LLM LLM p-value Non-LLM LLM p-value Perceived Reduction in Belief in Negative Thought LLM 0.26 $\pm$ 0.10 0.64 $\pm$ 0.13 0.02* 0.49 $\pm$ 0.12 0.65 $\pm$ 0.12 0.21 Perceived Reduction in Emotion Intensity 0.59 $\pm$ 0.13 0.89 $\pm$ 0.12 0.06 0.44 $\pm$ 0.10 0.65 $\pm$ 0.12 0.15 Perceived Relatability 5.16 $\pm$ 0.17 5.75 $\pm$ 0.13 ¡0.01** 4.48 $\pm$ 0.17 5.05 $\pm$ 0.18 0.03* Perceived Capacity of Inspiring Solutions 4.74 $\pm$ 0.18 5.46 $\pm$ 0.16 ¡0.01** 4.14 $\pm$ 0.19 4.76 $\pm$ 0.18 0.04* Likelihood of Future Engagement 4.53 $\pm$ 0.20 5.21 $\pm$ 0.17 ¡0.01** 3.92 $\pm$ 0.20 4.30 $\pm$ 0.20 0.15 Number of Words Written for Reflection 31.15 $\pm$ 2.28 44.22 $\pm$ 4.04 ¡0.01** 25.29 $\pm$ 2.42 33.56 $\pm$ 2.65 0.04*
@@ 4.2.1. Reevaluating Existing Thoughts and Beliefs
Participants who read LLM stories reported significantly greater perceived reduction in belief in their negative thought among crowdworkers ($Mean_{\text{LLM}}$ = 0.64 $\pm$ 0.13) compared to those in the Non-LLM condition ($Mean_{\text{Non-LLM}}$ = 0.26 $\pm$ 0.10, $p$ = 0.02). Among students, the difference followed the same direction ($Mean_{\text{LLM}}$ = 0.65 $\pm$ 0.12 vs. $Mean_{\text{Non-LLM}}$ = 0.49 $\pm$ 0.12), though it was not statistically significant ($p$ = 0.21).
@@
The reevaluation process also involved a reflective writing activity associated with the reflection question. As shown in Table 4, both crowdworkers and students wrote more words on average in response to the LLM stories. This activity aimed to provide participants with an opportunity to organize their thoughts and emotions into written form, which may have helped them restructure their thoughts to be more helpful. SL28 expressed one such change, saying:
@@
These findings suggest that both Non-LLM and LLM stories could support participants in reevaluating their thoughts and beliefs, offering them new insights and perspectives.
@@ 4.2.2. Potential of Stories in Inspiring Solutions and Actions
LLM stories received higher ratings for their ability to inspire potential solutions among both crowdworkers ($Mean_{\text{LLM}}$ = 5.46 $\pm$ 0.16) and students ($Mean_{\text{Non-LLM}}$ = 4.76 $\pm$ 0.18) compared to the Non-LLM condition ($Mean_{\text{LLM}}$ = 4.74 $\pm$ 0.18, $Mean_{\text{Non-LLM}}$ = 4.14 $\pm$ 0.19), with both differences reaching statistical significance ($p$ 0.01, $p$ = 0.04, respectively).
@@
Participants expressed that this reflection process involved more than just passive reading. The LLM stories helped participants to reflect on their personal struggles through the lens of the story and to actively engage with the suggested solutions. By presenting coping mechanisms embedded within the context of the narrative, participants felt the stories allowed them to see how similar strategies might apply in their own lives. The stories not only illustrated relatable challenges but also modeled potential solutions in a way that felt accessible and actionable. SL64 commented:
@@ 5. Discussion
Our study demonstrates how LLMs can significantly enhance human-generated content so that it is more relatable and impactful to new audiences. We found that these enhancements through dynamic personalization improved the utility of the stories as interventions for managing negative thoughts among young adults. These findings add to the growing body of literature on human-LLM collaboration (Chan et al., 2023; Denny et al., 2023; Bhattacharjee et al., 2024c; Guingrich and Graziano, 2023; Sharma et al., 2023a), which seeks to utilize and expand upon existing support mechanisms and content, while shining a light on the evaluation and scrutiny required to support the design of such tools for DMH (Bhattacharjee et al., 2024c; Korngiebel and Mooney, 2021; Demszky et al., 2023).
@@ 5.1. Key Insights
5.1.1. RQ1: Personalized LLM-Enhanced Stories in Exhibiting Narrative Qualities
@@
Our research highlights the promising capabilities of LLMs to dynamically generate content that can enhance the quality of narratives in DMH interventions. Similar to prior works (Sharma et al., 2023b; Kim et al., 2024; Bhattacharjee et al., 2024c), the LLM stories in our study processed open-ended inputs from participants and adapted the stories to their unique contexts. This capability underscores the potential of LLMs in crafting personalized stories that can support individuals in managing their specific barriers to psychological wellbeing.
@@
Our findings reveal that the LLM stories in our study enhanced most of the narrative qualities we investigated. The LLM stories were effective in communicating clear takeaways, which participants noted as essential for translating the stories’ lessons into their own lives (Bhattacharjee et al., 2022; Dyer, 1982). These stories were perceived to excel in promoting reflection, as explicit suggestions within the narratives encouraged readers to contemplate potential changes in their thinking patterns (Bhattacharjee et al., 2024a; Kocielnik et al., 2018). Additionally, the LLM stories balanced positivity with realistic portrayals of struggles. The stories often charted a structured progression from challenging life scenarios that resonated with participants’ experiences to outcomes showing small yet meaningful progress (Bhattacharjee et al., 2022).
@@
The LLM was able to render all of these benefits while still being comparable in authenticity to human-written stories, which likely reflects that the LLM effectively mirrored participants’ life experiences and that it drew on human-written example stories. However, while not statistically significant, we also note that the trend in the difference between the two conditions, favored the Non-LLM condition. Our qualitative data also revealed concerns about authenticity in the LLM condition, which could potentially reflect specific word choices made by the LLMs, participants’ familiarity with typical responses made by popular chatbots (Ariyaratne et al., 2023), and the occasional mechanical tone inherent in AI-generated content (Cheng et al., 2022). Recent studies have also examined text generated by ChatGPT across various contexts, noting an increased and repetitive occurrence of certain words such as ‘delve’, ‘intricate’, and ‘underscore’ (Kobak et al., 2024). The frequent use of certain relatively formal words or phrases may contribute to compromising the perceived authenticity of LLM-generated stories.
@@
5.1.2. RQ2: Personalized LLM-Enhanced Stories as an Intervention to Help People Manage Negative Thoughts